A Complete Bibliography of Applied Statistics: 1952–1959

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 4.03

Title word cross-reference

χ^{2} [199].

-Test [199].

10th [49]. 11th [254]. 1951 [4].

2 [187]. 2nd [312].

A. [76, 49, 254, 112, 363, 78]. Absence [120, 105]. Absence-Proneness [120]. Abuses [8]. Acceptance [303]. Accounting [195]. Accounts [113, 241, 364]. Accumulation [150]. Accuracy [31, 118, 9]. Achieve [261]. Ackoff [159]. Action [120, 130]. Active [136]. Adjusting [356]. Administration [81]. Advanced [326, 75]. Age [351]. Agricultural [359, 65, 77]. Aid [104]. Aitchison [269]. Alan [326]. Allen [253, 98]. Allocation [211]. Analysed [214]. Analysis [82, 229, 54, 48, 94, 140, 18, 327, 8, 247, 260, 306, 181, 328, 302, 296, 340, 262, 122, 145, 231, 160, 57, 285, 341, 152, 93, 324, 359, 92]. Answer [200, 142]. Answers [89, 184, 10, 34, 45]. any [197]. Application [348, 49, 254, 276, 138]. Applications [12, 24, 36, 47, 61, 73, 90, 110, 125, 143, 155, 173, 185, 201, 26, 312, 246, 132]. Applied [347, 140]. Appraisal [166]. Approach [295, 350, 42, 7, 253]. Aptitude [112]. Arbous [112]. Arithmetic [345]. Arthur [266]. Aspects [310, 315]. Assessments [353]. Associated [95]. Attributes [304]. Auction [277]. Auditing [350, 69]. Australia [354]. Authorities [113]. Automatic [214]. Average [261].

B [176, 204]. Back [99, 115, 133, 192, 207, 221, 365]. Bailey [325]. Bartlett [189]. Basic [304]. Basis [235]. Batteries [66]. be [84]. Beginners [238]. Bell [127]. Belts [9]. Bennett [160]. Berkeley [284]. Between [259, 149, 40, 211]. Bias [317]. Binomial [97]. Biological [332, 176, 77]. Biomathematics [176]. Biometric [75]. Biometrika [146]. Bishop [359]. Block [306]. Blocks [198]. Blood [274]. Blood-Clotting [274]. Boddington [49]. Book [76, 288, 218, 310, 163, 203, 97, 220, 206, 237, 256, 238, 270, 286, 326, 348, 324, 362, 282, 283, 361, 49, 330, 254, 269, 347, 345, 243, 26, 75, 48, 94, 130, 146, 50, 112, 323, 327, 363, 312, 176, 127, 328, 302, 27, 79, 95, 96, 157, 189, 188, 267, 340, 114, 346, 359, 113, 159, 98, 241, 299, 268, 325, 187, 145, 364, 271, 242, 342, 343, 236, 266, 190, 235, 287, 300, 360, 132, 128, 77, 158, 162, 161, 240, 253, 78, 204, 51, 92, 160, 284, 311, 285, 14, 129, 205, 239, 341, 344, 93]. Borstal [218]. Both [250]. Branch [21, 170]. Britain [212, 55]. British [121]. Broadcast [213]. Brown [269]. Buckingham [287]. Buckland [268]. Budget [336]. Budgets [364]. Bush [203]. Business [364].

C [288, 220, 269, 26, 302, 157, 364, 51]. C. [75, 359, 92, 239]. Calculating [232]. Calculation [150]. Calculations [79]. Calibration [20]. Calvin [158]. Cardpack [187]. Cards [335]. Care [117]. Carl [160]. Carlo [210]. Cauter [205]. Cave [128]. Cedric [176]. Census [39, 271]. Censuses [212]. Certain [261]. Chacon [237]. Chance [187]. Change [44, 71]. Chart [278, 320, 180]. Charts [66, 258]. Chemical [270, 18, 42, 160]. Chemistry [160]. Chemists [27]. Chessboard [187]. Child [363]. Choice [319, 187]. Cigarette [334]. Circle [63]. Class [248, 5]. Classification [352, 332, 241]. Claus [344]. Clotting [274]. Coal [31]. Coding [241]. Coke [355]. Collaborative [18]. Comment [215]. Comments [11, 23, 35, 46, 60, 72, 88, 109, 124, 154, 172, 183, 216, 233, 251, 264, 280, 297, 308, 321, 357]. Commerce [49, 254]. Communication [205]. Companies [351]. Comparison [139, 43]. Compilation [190]. Complex [140, 181]. Composite [33]. Computation [335]. Computers [214, 246, 337]. Concise [163]. Conditions [171, 261]. Confidence [263]. Congress [4]. Connection [195]. Construction [215, 107, 122, 86]. Consumer [318, 30]. Consumption [41]. Continuous [262]. Contrôle [128]. Contributions [204]. Control [66, 230, 101, 179, 278, 258, 305, 315, 180, 161, 64, 311]. Controlled [171]. Conversion [319]. Conveyor [9]. Copybook [238]. Cornell [242]. Correlated [220]. Correlation [54, 324]. Correlations [197]. Cost [168, 278, 81, 294]. Costs [354, 302]. Cotton [315, 70]. Course [342, 341]. Cowden [311]. Cox [362]. Crisi [307]. Critique [162]. Current [165, 162]. Curso [237]. Cybernetics [361].

D [359, 98]. D. [163, 362, 127, 235]. Dahlberg [266]. Daily [105]. Data [276, 181, 317]. Davies [270, 346, 145]. Decision [282]. Decontamination [198]. Defectives [58, 119]. Deferred [333]. Degrees [33]. Demand [328, 102]. Derivation [319]. Descriptiva [237]. Design [209, 171, 6, 48, 94, 306, 249, 159, 145, 235, 93]. Designing [336]. Detect [231]. Detergent [68]. Determination [119, 170]. Development [57, 93]. Deviates [220].

Dictionary [268]. Difference [151]. Differences [259, 40]. Different [226]. Digit [43]. Discontinued [319]. Discriminant [248]. Dishwashing [68]. Distribution [326, 39, 269, 305, 228]. Domestic [68]. Donald [93]. Doris [256]. Dornbusch [239]. Downham [205]. DuBois [324]. Dudley [311]. Dynamic [240].

E. [220, 206, 146, 236]. Econometrics [225, 103, 284]. Economic [345, 248, 359, 275, 55]. Economics [58, 269, 204]. Economist [292]. Economists [347]. Economy [121]. Edey [364]. Edge [335]. Edge-Punched [335]. Edition [312, 302, 270, 49, 254]. Editor [123, 108, 153, 338, 59, 87, 141]. Editorial [2, 16, 291]. Educational [242]. Edward [98]. Effect [226]. Effects [213]. Eight [198]. Elasticities [102]. Electrical [127]. Electronic [246]. Elementary [256]. Elements [157, 337]. Elimination [317]. Ely [98]. Empirical [107]. Engineering [76, 206, 249, 127]. Enrick [161]. Enrique [237]. Environmental [171]. Epidemics [325]. Epidemiology [360]. Error [232]. Errors [106, 231]. Essentials [242]. Establishments [170]. Estadistica [237]. Estimates [120]. Estimating [17, 40, 32, 5]. Evaluation [167, 68]. Evolutionary [245]. Examination [317]. Examining [181]. Examples [194, 76]. Executive [120, 137]. Exercises [114]. Expenditure [182, 32]. Experience [30]. Experienced [259]. Experiences [54]. Experiment [198, 196, 122, 65, 70, 94]. Experimental [283, 140, 316, 235]. Experiments [229, 214, 104, 93, 362, 48, 145]. Exports [67]. Extension [199].

F [327, 158, 239]. F. [203]. Fabrications [128]. Factor [319, 8, 181, 296, 57]. Factorial [229, 198, 122]. Factors [240].

Factory [302]. Facts [14]. Family [226, 32]. Fascination [286]. Feller [312]. Field [182]. Fieller [220]. Figures [137, 14]. Finite [356]. Finney [235]. Firm [81]. First [271, 342]. Fisher [188, 267, 77]. Flowers [18]. Flowmeter [20]. Fluctuations [41]. Foodstuffs [166, 181]. Forecasting [167, 55]. Foreword [223]. Formulas [256, 78]. Forward [1]. Foundations [307]. Four [198]. Francis [242]. Frank [77]. Franklin [160]. Frederick [327]. Freedom [33]. Front [15, 52, 100, 116, 134, 193, 208, 222]. Fryer [157]. Functions [248, 121]. Fundamental [166]. Fundamentals [330]. Funds [178]. Future [17].

G [97, 206, 326, 130, 50, 112, 114, 113, 98, 268, 242, 343, 341]. G. [361]. Games [243]. Gas [20]. George [204]. Gets [83]. Glass [29]. Good [84]. Goulden [92]. Grant [206]. Graphic [158]. Graphical [151, 122]. Graphing [137]. Great [212, 55]. Grenander [285]. Ground [18]. Group [247, 260]. Guide [256, 197, 236, 303]. Guiding [120]. Guilbaud [361].

H

[324, 330, 347, 26, 94, 302, 95, 340, 190, 51, 92]. H. [146, 157, 360, 129]. Hald [76, 78]. Handbook [206, 158]. Harold [364]. Harry [97, 323, 253]. Hartley [146]. Henry [323]. Hermann [218]. Hilde [363]. Himmelweit [363]. History [163]. Hogben [307, 187]. Hospital [209]. Household [182]. Hugh [190].

Ideas [205]. II [260, 122]. Ilersic [49, 254]. Income [182, 165, 266]. Increasing [245]. Index [19, 215, 319, 136, 67, 86]. Individual [32]. Industrial [12, 24, 36, 47, 61, 73, 90, 110, 125, 143, 155, 173, 185, 201, 352, 206, 282, 245, 210, 139, 121, 212, 145, 240, 284]. Industry [270, 3, 160]. Infection [148].

Inference [288, 267, 129]. Information [38, 53, 80, 135, 147, 177, 224, 244, 273, 290, 314, 331, 366, 226]. Informats [226]. Initial [279]. Inquiry [226]. Inspection [304]. Integrating [20]. Intermediate [231]. International [98, 4]. Interview [250]. Interviewer [194]. Interviewers [259]. Introduction [288, 3, 312, 79, 189, 359, 300]. Inventories [328]. Investigation [18, 106, 344]. Investment [178]. Ireson [206].

J [163, 286, 327, 27, 325, 235, 300, 311, 14]. J. [310, 283, 269, 79, 113, 98, 241, 360, 284, 129, 205]. James [204]. Jerome [288]. John [96, 204]. Johnson [256]. Jurors [71]. Jury [6].

Karmel [347]. Katona [204]. Kempthorne [48]. Kendall [326, 50, 114, 268, 341]. Kingdom [50]. Klein [204].

L [206, 270, 159, 145, 160]. L. [26, 51]. Lancelot [187]. Landblom [238, 238]. Lansing [204]. Lawrence [204]. Learning [203]. Length [334]. Leonard [302]. Leslie [218]. Lester [49]. Letter [108, 153, 338, 59, 87, 141]. Letters [123, 279]. Lev [129]. Level [168]. Lewis [220, 162]. Li [288]. Limitations [225]. Limits [179, 263]. Line [22]. Linear [243, 227, 138]. Lives [29]. Lloyd [161]. Local [113]. Lognormal [269, 305]. Looks [4]. Lot [356]. Loveday [342]. Lyle [302].

M [328, 241, 129, 239]. M. [330, 94, 50, 95, 189, 340, 114, 285, 14]. Madge [96]. Magazines [169]. Maintenance [328]. Makeshift [140]. Making [29]. Management [4, 206]. Mannheim [218]. Manufacture [168, 42]. Marginal [294]. Market [299]. Marris [345]. Martin [310]. Massed [66]. Matching [332]. Matematico [237].

Material [316]. Materials [231, 138, 211]. Mathematical [325, 284]. Mathematics [163, 130, 176]. Matter [15, 52, 99, 115, 133, 100, 116, 134, 192, 207, 221, 193, 208, 222, 365]. Maurice [326, 268, 341]. Maverick [162]. Max [346]. McCarthy [327, 300]. Mean [33]. Means [139]. Measured [149]. Measurement [102, 105, 21, 231]. Measurements [262, 95]. Measuring [169]. Medical [63, 28, 117, 77]. Medicine [352, 83]. Meetings [217, 234, 252, 265, 281, 298, 309, 322, 339, 358]. Melman [240]. Method [245, 137]. Methods [218, 270, 75, 140, 3, 18, 127, 210, 27, 188, 267, 204, 51, 92, 311, 344, 287, 161]. Michel [346]. Mill [70]. Minimisation [295]. Minimum [168]. Missing [306, 214]. Modelo [237]. Models [203]. Modified [179]. Monte [210]. Monthly [355]. Moore [343]. Morbidity [226]. Morgan [204]. Moroney [14]. Morris [256, 282]. Morse [328]. Mortality [293]. Moser [344]. Mosteller [203]. Motivation [323]. Mounsey [79]. Multi [231]. Multi-Stage [231]. Multiple [260, 149, 340, 170]. Multivariate [324, 341, 152].

N [256, 363, 204]. N. [325]. National [165, 266]. Nature [50]. Neale [236]. Neglected [86]. Nellie [238]. Net [67]. Newspapers [169]. Neyman [284]. Nomogram [85]. Nomograms [107, 122]. Non [151]. Non-Parametric [151]. Norbert [161]. Normal [220]. Normality [258]. Norman [160]. Notes [11, 23, 35, 46, 60, 72, 88, 109, 124, 154, 172, 183, 216, 233, 251, 264, 280, 297, 308, 321, 357]. Number [43]. Numbers [19, 215, 86, 286]. Numerical [287].

O [146, 266]. O. [270, 130, 145]. Observation [106]. Observational [276]. Observations [306]. Obtaining [226]. Official [275, 236]. One [316, 22].

Operation [245]. Operational [328, 346]. Operations [302]. Opinions [327]. Oppenheim [363]. Optical [262]. Organisation [170]. Orgin [57]. Oscar [48]. Other [131].

P [310, 236]. P. [347, 343]. Pamela [363]. Pandurang [132]. Panel [30]. Paper [348]. Parametric [151]. Part [247, 260, 107, 122]. Pattern [318]. Pearson [220, 146]. Pemberton [360]. Percentage [119]. Percentages [320]. Performance [68, 21]. Periodograph [262]. Personnel [180]. Petroleum [41]. Philip [324, 327, 328, 302, 300]. Pilot [271]. Plane [149]. Planning [209, 362, 70]. Plans [356]. Plastics [7]. Point [86]. Points [149]. Population [17, 271, 190]. Postal [250]. Pots [29]. Practical [167]. Practice [227, 346]. Prediction [218, 332, 56]. Predictive [197]. Preliminary [317]. Prescribing [310]. Presentation [229, 283, 158]. Price [215, 86]. Prices [19]. Primer [239]. Principle [332]. Principles [304, 176, 343]. Probability [312, 284]. Probits [276]. Problem [352, 230, 42]. Problems [166, 195, 210]. Procedure [261, 327]. Procedures [58]. Proceedings [284]. Process [231]. Processes [44, 282, 189]. Production [270, 230, 302, 121, 81]. Productivity [136, 245, 162, 240]. Products [211]. Professor [307]. Programming [227, 138, 243]. Proneness [120]. Proportion [232, 71]. Pseudo [337]. Pseudo-Random [337]. Psychometry [284]. Public [351]. Publication [272]. Publications [131, 164, 175, 191, 219, 255, 289, 301, 313, 329, 349]. Pulmonary [148]. Punched [335]. Purchases [318]. Pyrethrum [18].

Qualification [6]. Quality [101, 168, 258, 305, 315, 161, 64, 311, 355].

Quarterly [355]. Quenouille

[330, 94, 95, 340]. Question [200, 142]. Questions [89, 184, 10, 34, 45]. Queueing [117]. Queues [328]. Quoted [351].

R [288, 203, 362, 49, 254, 268, 204]. R. [203, 159, 98, 342, 287, 128]. Radar [263]. Radhakrishna [75]. Radioactive [198]. Random [220, 104, 337, 43]. Randomised [306]. Range [263]. Rao [75]. Rates [293]. Rationalization [282]. Raw [138]. Readership [169]. Readings [40]. Reasoning [330, 300]. Received [131, 164, 175, 191, 219, 255, 272, 289, 301, 313, 329, 349]. Records [195]. Reduction [283]. Reference [270, 269]. Regression [247, 260, 302, 149, 278, 180, 231]. Regulators [101]. Reichmann [286]. Relation [218, 81]. Reliability [353]. Research [238, 270, 188, 346, 284, 93, 75, 323, 159, 299, 77]. Reservoirs [148]. Results [82, 229, 140, 167, 122, 283]. Retail [19]. Returns [65]. Review [176, 14]. Reviews [76, 288, 218, 310, 163, 203, 97, 220, 206, 237, 256, 238, 270, 286, 326, 348, 324, 362, 282, 283, 361, 49, 330, 254, 269, 347, 345, 243, 26, 75, 48, 94, 130, 146, 50, 112, 323, 327, 363, 312, 127, 328, 302, 27, 79, 95, 96, 157, 189, 188, 267, 340, 114, 346, 359, 113, 159, 98, 241, 299, 268, 325, 187, 145, 364, 271, 242, 342, 343, 236, 266, 190, 235, 287, 300, 360, 132, 128, 77, 158, 162, 161, 240, 253, 78, 204, 51, 92, 160, 284, 311, 285, 129, 205, 239, 341, 344, 93]. Richardson [283]. Risk [241]. Road [257]. Robert [256]. Roberts [253]. Robin [345]. Romig [97]. Ronald [188, 267, 77]. Rosenblatt [285]. Royal [12, 24, 36, 13, 25, 37, 47, 61, 73, 63, 62, 74, 90, 110, 125, 91, 111, 126, 143, 155, 173, 144, 156, 174, 185, 201, 217, 186, 202, 234, 252, 265, 281, 298, 309, 322, 339, 358].

S [220, 146, 189, 241, 205]. S. [243, 239].

Sale [277]. Sales [167, 21, 137]. Same [316]. Sample [182, 6, 194, 250, 119]. Samples [69]. Sampling [58, 327, 304, 356, 31, 195, 9, 65, 43, 132, 257, 303, 228]. Scarce [211]. Schemes [333, 119]. Schmid [158, 239]. Science [176, 96]. Scientific [267]. Scope [28, 225]. Search [263]. Section [12, 24, 36, 13, 25, 37, 47, 61, 73, 62, 74, 90, 110, 125, 91, 111, 126, 143, 155, 173, 144, 156, 174, 185, 201, 186, 202]. Sections [217, 234, 252, 265, 281, 298, 309, 322, 339, 358]. Selection [180, 138]. Selector [104]. Sensory [353]. Sentencing [333]. Sequential [58, 261]. Serial [43]. Series [319, 139, 340, 285]. Service [136]. Set [316]. Setting [101]. Several [171]. Seymour [240]. Sheet [149]. Shipping [354]. Shop [170]. Short [41]. Short-Term [41]. Significance [151, 320]. Simple [295, 151, 293, 247, 261]. Single [356, 119]. Single-Sample [119]. Size [295, 356, 81]. Skew [228]. Smith [176]. Social [310, 96, 113, 159, 5, 239, 344]. Societ [234]. Society [12, 24, 36, 13, 25, 37, 47, 61, 73, 63, 62, 74, 90, 110, 125, 91, 111, 126, 143, 155, 173, 144, 156, 174, 185, 201, 217, 186, 202, 252, 265, 281, 298, 309, 322, 339, 358]. Socio [248]. Socio-Economic [248]. Some [85, 140, 304, 195, 246, 167, 315]. Sources [50, 299]. Special [270, 269]. Specification [168, 7]. Spinning [70]. Squares [33]. Staff [170]. Stage [250, 231]. Standard [232, 261]. Standardising [66]. States [212]. Stationary [285]. Statistical [12, 24, 36, 13, 25, 37, 76, 47, 61, 73, 63, 62, 74, 90, 110, 125, 91, 111, 126, 143, 155, 173, 144, 156, 174, 185, 201, 217, 186, 202, 234, 252, 265, 281, 298, 309, 322, 339, 358, 288, 256, 270, 330, 75, 3, 127, 350, 27, 79, 188, 267, 246, 42, 299, 268, 343, 235, 300, 77, 161, 78, 92, 160, 311, 7, 285, 335, 129, 93]. Statistician [4]. Statisticians [146]. Statistics [63, 28, 209, 44, 326, 348, 49, 85, 254, 347, 50, 277, 178, 56, 118, 104, 307, 107, 122, 275, 342, 165, 190, 292, 253, 284, 239, 26, 157, 114, 98, 242, 236, 51]. Statistique [128]. Statler [93]. Stephan [327]. Stochastic [44, 203, 189]. Store [295]. Store-Size [295]. Stores [350]. Structure [351]. Struik [163]. Stuart [326]. Stubs [334]. Students [176]. Studies [274, 360]. Study [13, 25, 37, 63, 62, 74, 91, 111, 126, 144, 156, 174, 186, 202, 198, 353, 29, 249]. Studying [213]. Subjective [352, 166]. Successive [40]. Sudan [271]. Sukhatme [132]. Supply [328]. Surnames [279]. Survey [327, 250, 336, 307, 204, 344]. Surveys [182, 194, 132]. Sutton [130]. Symposium [284]. Systematic [9, 43]. Systems [328].

T [218, 238, 282, 283, 361, 363, 325]. T. [220, 205]. Tables [146, 112, 77, 78, 97]. Taken [194]. Tasting [196]. Teaching [104]. Technique [66, 213, 181]. Techniques [343]. Technological [26]. Television [213, 363]. Term [41]. Terms [354, 268]. Test [274, 82, 151, 199, 261]. Testers [112]. Testing [171, 348]. Textile [82]. Textiles [315]. Their [49, 254, 33, 190]. Theoretical [114]. Theory [76, 326, 243, 312, 325, 132, 326]. Third [270, 302, 284]. Three [247, 260]. Three-Group [247, 260]. Time [139, 285, 340]. Time-Series [340]. Tippett [26, 302, 51]. Tolerances [355]. Too [84]. Tools [96]. Totals [32]. Toussaint [359]. Towns [5]. Trade [354, 98]. Traffic [257]. Training [218]. Treatments [171, 316]. Trends [275, 165]. True [107, 84]. Tuberculosis [148]. Two [274, 194, 22].

U. [285]. United [50, 212]. USA [64].

Usage [162]. Use [120, 6, 316, 178, 262, 199, 107, 122, 180, 231, 137, 335, 69]. Used [29]. User [30]. Uses [85, 269, 8]. Utting [113].

V [132, 253, 284]. Vajda [243]. Value [197]. Values [214]. Variability [274, 194, 40, 249].

Variable [328]. Variables [149, 228]. Variance [306]. Variations [105]. Verhulst [346]. Very [228]. Villars [93]. Vince [363]. Visual [82]. Visualized [266]. Vol [284]. Voltage [101]. Volume [38, 53, 80, 135, 147, 177, 224, 244, 273, 290, 314, 331, 366, 237, 326, 67, 146, 50, 312, 187].

W. [206, 286, 27, 359, 253]. Walker [129]. Wallis [253]. Warwickshire [148]. Weekly [355]. Whether [261]. Who [83]. Wilkins [218]. Willard [360]. William [282, 312, 268]. Wolfenden [190]. Work [182, 238]. Workers [188].

Yarn [82]. Yates [77]. Youden [27].

Zealand [354, 236].

References

Anonymous:1952:F

[1] Anonymous. Forward. Applied Statistics, 1(1):1, March 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:Ea

[2] Anonymous. Editorial. Applied Statistics, 1(1):2, March 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dudding:1952:ISM

[3] Bernard P. Dudding. The introduction of statistical methods to industry. Applied Statistics, 1(1):3–20, March 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swan:1952:SLI

[4] Albert W. Swan. A statistician looks at The International Management Congress of 1951. Applied Statistics, 1 (1):21–26, March 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkins:1952:ESC

[5] Leslie T. Wilkins. Estimating the social class of towns. Applied Statistics, 1(1): 27–33, March 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1952:UJQ

[6] Thomas Corlett. A use for the jury qualification in sample design. Applied Statistics, 1(1):34–36, March 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wainwright:1952:SAS

[7] Charles Wainwright. A statistical approach to the specification of plastics. Applied Statistics, 1(1):37–44, March 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Eysenck:1952:UAF

[8] Hans J. Eysenck. Uses and abuses of factor analysis. Applied Statistics, 1(1): 45–49, March 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1952:ASS

[9] Geoffrey H. Jowett. The accuracy of systematic sampling from conveyor belts. Applied Statistics, 1(1):50–59, March 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Salter:1952:QA

[10] G. Salter, L. T. Wilkins, and G. H. Jowett. Questions and answers. Applied Statistics, 1(1):60–62, March 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:NCa

[11] Anonymous. Notes and comments. Applied Statistics, 1(1):63–64, March 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:IASa

[12] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 1(1):65–73, March 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:SSRa

[13] Anonymous. Study section of the royal statistical society. Applied Statistics, 1 (1):74–79, March 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkins:1952:BRF

[14] L. T. Wilkins. Book review: Facts from Figures, by M. J. Moroney. Applied Statistics, 1(1):80, March 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:FMP

[15] Anonymous. Front matter. Applied Statistics, 1(1):[3], March 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:Eb

[16] Anonymous. Editorial. Applied Statistics, 1(2):81, June 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1952:EFP

[17] Peter R. Cox. Estimating the future population. Applied Statistics, 1(2): 82–94, June 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dyke:1952:CIM

[18] George V. Dyke and Emily P. Simpson. A collaborative investigation into methods of chemical analysis of ground pyrethrum flowers. Applied Statistics, 1(2):95–105, June 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Allen:1952:INR

[19] Roy G. D. Allen. Index numbers of retail prices 1938–51. Applied Statistics, 1(2): 106–117, June 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Spendley:1952:CIG

[20] William Spendley. Calibration of an integrating gas flowmeter. Applied Statistics, 1(2):118–125, June 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ramm:1952:MBS

[21] Bernard T. Ramm. The measurement of branch sales performance. Applied Statistics, 1(2):126–130, June 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jessop:1952:OLT

[22] William N. Jessop. One line or two? Applied Statistics, 1(2):131–137, June 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:NCb

[23] Anonymous. Notes and comments. Applied Statistics, 1(2):138–139, June 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:IASb

[24] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 1(2):140–144, June 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:SSRb

[25] Anonymous. Study section of the royal statistical society. Applied Statistics, 1 (2):145–147, June 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1952:BRT

[26] O. L. Davies. Book reviews: Technological Applications of Statistics, by L. H. C. Tippett. Applied Statistics, 1 (2):148, June 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1952:BRS

[27] G. H. Jowett. Book reviews: Statistical Methods for Chemists, by W. J. Youden. Applied Statistics, 1(2):148, June 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1952:SMS

[28] Norman T. J. Bailey. The scope of medical statistics. Applied Statistics, 1(3): 149–162, November 1952. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Keen:1952:SLP

[29] Joan Keen and Denys J. Page. A study of the lives of pots used in making glass. Applied Statistics, 1(3):163–168, November 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wadsworth:1952:EUC

[30] Robert N. Wadsworth. The experience of a user of a consumer panel. Applied Statistics, 1(3):169–178, November 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hebden:1952:ASC

[31] Jack Hebden and Geoffrey H. Jowett. The accuracy of sampling coal. Applied Statistics, 1(3):179–191, November 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemsley:1952:EIE

[32] W. F. F. Kemsley. Estimating individual expenditure from family totals. Applied Statistics, 1(3):192–201, November 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Johnson:1952:CMS

[33] Leroy P. V. Johnson and Ernest S. Keeping. Composite mean squares and their degrees of freedom. Applied Statistics, 1(3):202–205, November 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shankleman:1952:QA

[34] E. Shankleman and O. L. Davies. Questions and answers. Applied Statistics, 1(3):206–207, November 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:NCc

[35] Anonymous. Notes and comments. Applied Statistics, 1(3):208–209, November 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:IASc

[36] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 1(3):210, November 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:SSRc

[37] Anonymous. Study section of the royal statistical society. Applied Statistics, 1(3):211–212, November 1952. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1952:VIP

[38] Anonymous. Volume information. Applied Statistics, 1(3):[2], November 1952. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clark:1953:CD

[39] Dudley A. Clark. The census of distribution. Applied Statistics, 2(1):1– 12, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Keen:1953:EVD

[40] Joan Keen and Denys J. Page. Estimating variability from the differences between successive readings. Applied Statistics, 2(1):13–23, March 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Butler:1953:STF

[41] John Douglas Butler. Short-term fluctuations in petroleum consumption. Applied Statistics, 2(1):24–29, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Muir:1953:SAP

[42] Andrew Muir and Frank Downton. A statistical approach to a problem of chemical manufacture. Applied Statistics, 2(1):30–38, March 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Price:1953:CSN

[43] Orville O. Price and Alfred J. McKeon. A comparison of serial number digit sampling with systematic and random sampling. Applied Statistics, 2(1): 39–43, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartlett:1953:SPS

[44] Maurice S. Bartlett. Stochastic processes or the statistics of change. Applied Statistics, 2(1):44–64, March 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanRest:1953:QA

[45] E. D. Van Rest and W. J. Martin. Questions and answers. Applied Statistics, 2 (1):65–66, March 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:NCa

[46] Anonymous. Notes and comments. Applied Statistics, 2(1):67, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:IASa

[47] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 2(1):68–69, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1953:BRD

[48] O. L. Davies. Book reviews: The Design and Analysis of Experiments, by Oscar Kempthorne. Applied Statistics, 2 (1):70, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Booker:1953:BRS

[49] H. S. Booker. Book reviews: Statistics and Their Application to Commerce. 10th Edition, by A. Lester Boddington and A. R. Ilersic. Applied Statistics, 2 (1):70–71, March 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Devons:1953:BRS

[50] E. Devons. Book reviews: The Sources and Nature of the Statistics of the United Kingdom. Volume I, by M. G. Kendall. Applied Statistics, 2(1):71–72, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanRest:1953:BRM

[51] E. D. Van Rest. Book reviews: The Methods of Statistics, by L. H. C. Tippett. Applied Statistics, 2(1):72, March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:FMP

[52] Anonymous. Front matter. Applied Statistics, 2(1):[2], March 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:VIPa

[53] Anonymous. Volume information. Applied Statistics, 2(1):[2], March 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1953:ECA

[54] Donald G. Beech. Experiences of correlation analysis. Applied Statistics, 2(2): 73–85, June 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shankleman:1953:EFG

[55] Eric Shankleman. Economic forecasting in Great Britain. Applied Statistics, 2 (2):86–100, June 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hughes:1953:SP

[56] Cyril Davenport Hughes. Statistics and prediction. Applied Statistics, 2(2):101– 106, June 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vincent:1953:ODF

[57] Douglas F. Vincent. The orgin and development of factor analysis. Applied Statistics, 2(2):107–117, June 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Champernowne:1953:ESS

[58] David G. Champernowne. The economics of sequential sampling procedures for defectives. Applied Statistics, 2(2):118–130, June 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nicholson:1953:LE

[59] J. L. Nicholson and W. F. F. Kemsley. Letter to the editor. Applied Statistics, 2(2):131–132, June 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:NCb

[60] Anonymous. Notes and comments. Applied Statistics, 2(2):133–135, June 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:IASb

[61] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 2(2):136–137, June 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:SSRa

[62] Anonymous. Study section of the royal statistical society. Applied Statistics, 2 (2):138–139, June 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:SCM

[63] Anonymous. Study circle on medical statistics: Royal statistical society. Applied Statistics, 2(2):140, June 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanRest:1953:QCU

[64] Edward D. Van Rest. Quality control in the USA. Applied Statistics, 2(3): 141–151, November 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Palca:1953:ESA

[65] Henry Palca. An experiment in the sampling of agricultural returns. Applied Statistics, 2(3):152–159, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beer:1953:TSM

[66] Stafford Beer. A technique for standardising massed batteries of control charts. Applied Statistics, 2(3):160–165, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Billewicz:1953:IVN

[67] Wladyslaw Z. Billewicz. An index of the volume of net exports. Applied Statistics, 2(3):166–171, November 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pugh:1953:EDP

[68] Clifford Pugh. The evaluation of detergent performance in domestic dishwashing. Applied Statistics, 2(3):172–179, November 1953. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1953:USA

[69] Noel Williams. The use of samples in auditing. Applied Statistics, 2(3):180–183, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peake:1953:PEC

[70] Robert E. Peake. Planning an experiment in a cotton spinning mill. Applied Statistics, 2(3):184–192, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1953:CPJ

[71] Thomas Corlett. A change in the proportion of jurors. Applied Statistics, 2(3):193–195, November 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:NCc

[72] Anonymous. Notes and comments. Applied Statistics, 2(3):196, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:IASc

[73] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 2(3):197–199, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:SSRb

[74] Anonymous. Study section of the royal statistical society. Applied Statistics, 2(3):200–201, November 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1953:BRA

[75] O. L. Davies. Book reviews: Advanced Statistical Methods in Biometric Research, by C. Radhakrishna Rao. Applied Statistics, 2(3):202, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:BRS

[76] Anonymous. Book reviews: Statistical Theory with Engineering Examples, by A. Hald. Applied Statistics, 2(3):203, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1953:BRS

[77] L. H. C. Tippett. Book reviews: Statistical Tables for Biological, Agricultural and Medical Research, by Ronald A. Fisher and Frank Yates. Applied Statistics, 2(3):203, November 1953. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Turner:1953:BRS

[78] F. C. Turner. Book reviews: Statistical Tables and Formulas, by A. Hald. Applied Statistics, 2(3):203, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1953:BRI

[79] G. H. Jowett. Book reviews: An Introduction to Statistical Calculations, by J. Mounsey. Applied Statistics, 2(3):204, November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1953:VIPb

[80] Anonymous. Volume information. Applied Statistics, 2(3):[2], November 1953. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Melman:1954:PAC

[81] Seymour Melman. Production and administration cost in relation to size of firm. Applied Statistics, 3(1):1– 11, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baker:1954:ARV

[82] Anthony G. Baker, Kenneth W. Hillier, and Robert H. Woodward. An analysis of the results of a visual test of textile yarn. Applied Statistics, 3(1):12– 18, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gray:1954:WGM

[83] Percy G. Gray and Ann Cartwright. Who gets the medicine? Applied Statistics, 3(1):19–28, March 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stuart:1954:TGT

[84] Alan Stuart. Too good to be true? Applied Statistics, 3(1):29–32, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Broadbent:1954:SUN

[85] Simon Broadbent. Some uses of the nomogram in statistics. Applied Statistics, 3(1):33–43, March 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ruderman:1954:NPC

[86] Armand Peter Ruderman. A neglected point in the construction of price index numbers. Applied Statistics, 3(1): 44–47, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Robson:1954:LE

[87] Peter Robson. Letter to the editor. Applied Statistics, 3(1):48–49, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:NCa

[88] Anonymous. Notes and comments. Applied Statistics, 3(1):50–51, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lyle:1954:QA

[89] Philip Lyle and E. Shankleman. Questions and answers. Applied Statistics, 3 (1):52–54, March 1954. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:IASa

[90] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 3(1):55–57, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:SSRa

[91] Anonymous. Study section of the royal statistical society. Applied Statistics, 3 (1):58, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanRest:1954:BRM

[92] E. D. Van Rest. Book reviews: Methods of Statistical Analysis, by C. H. Goulden. Applied Statistics, 3(1): 59, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vanRest:1954:BRS

[93] E. D. van Rest. Book reviews: Statistical Design and Analysis of Experiments for Development Research, by Donald Statler Villars. Applied Statistics, 3(1): 59–60, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1954:BRD

[94] O. L. Davies. Book reviews: The Design and Analysis of Experiment, by M. H. Quenouille. Applied Statistics, 3(1): 60–61, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1954:BRA

[95] G. H. Jowett. Book reviews: Associated Measurements, by M. H. Quenouille. Applied Statistics, 3(1):61–62, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1954:BRT

[96] G. H. Jowett and M. Hodges. Book reviews: The Tools of Social Science, by John Madge. Applied Statistics, 3(1): 62, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1954:BRB

[97] D. G. Beech. Book reviews: 50–100 Binomial Tables, by Harry G. Romig. Applied Statistics, 3(1):63, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nuttall:1954:BRI

[98] C. Nuttall. Book reviews: International Trade Statistics, by R. G. D. Allen and J. Edward Ely. Applied Statistics, 3(1): 63–64, March 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:BMPa

[99] Anonymous. Back matter. Applied Statistics, 3(1):[1], March 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:FMPa

[100] Anonymous. Front matter. Applied Statistics, 3(1):[2], March 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Desmond:1954:QCS

[101] David J. Desmond. Quality control on the setting of voltage regulators. Applied Statistics, 3(2):65–73, June 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leser:1954:MED

[102] Conrad E. V. Leser. The measurement of elasticities of demand. Applied Statistics, 3(2):74–84, June 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shankleman:1954:WE

[103] Eric Shankleman. What is econometrics? Applied Statistics, 3(2):85– 89, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Knowles:1954:ERS

[104] Elsie A. G. Knowles. Experiments with a random selector as an aid to the teaching of statistics. Applied Statistics, 3(2): 90–103, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liddell:1954:MDV

[105] F. D. K. Liddell. The measurement of daily variations in absence. Applied Statistics, 3(2):104–111, June 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leslie:1954:IEO

[106] Rupert Thomas Leslie. Investigation of errors of observation. Applied Statistics, 3(2):112–115, June 1954. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lyle:1954:CNUa

[107] Philip Lyle. The construction of nomograms for use in statistics: Part I. true and empirical nomograms. Applied Statistics, 3(2):116–125, June 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fowler:1954:LE

[108] R. F. Fowler. Letter to the editor. Applied Statistics, 3(2):126–127, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:NCb

[109] Anonymous. Notes and comments. Applied Statistics, 3(2):128–129, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:IASb

[110] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 3(2):130–132, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:SSRb

[111] Anonymous. Study section of the royal statistical society. Applied Statistics, 3 (2):133, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Elliott:1954:BRT

[112] E. Elliott. Book reviews: Tables for Aptitude Testers, by A. G. Arbous. Applied Statistics, 3(2):134, June 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moser:1954:BRS

[113] C. A. Moser. Book reviews: Social Accounts of Local Authorities, by J. E. G. Utting. Applied Statistics, 3(2):134– 135, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Knowles:1954:BRE

[114] E. A. G. Knowles. Book reviews: Exercises in Theoretical Statistics, by M. G. Kendall. Applied Statistics, 3(2): 135–136, June 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:BMPb

[115] Anonymous. Back matter. Applied Statistics, 3(2):[2], June 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:FMPb

[116] Anonymous. Front matter. Applied Statistics, 3(2):[2], June 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1954:QMC

[117] Norman T. J. Bailey. Queueing for medical care. Applied Statistics, 3(3): 137–145, November 1954. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hughes:1954:AS

[118] Cyril Davenport Hughes. Accuracy in statistics. Applied Statistics, 3(3):146– 149, November 1954. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Horsnell:1954:DSS

[119] Gareth Horsnell. The determination of single-sample schemes for percentage defectives. Applied Statistics, 3(3):150– 158, November 1954. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arbous:1954:UEA

[120] Adrian Garth Arbous and Herbert Simon Sichel. The use of estimates of absence-proneness for guiding executive action. Applied Statistics, 3(3):159–173, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leser:1954:PFB

[121] Conrad E. V. Leser. Production functions for the British industrial economy. Applied Statistics, 3(3):174–183, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lyle:1954:CNUb

[122] Philip Lyle. The construction of nomograms for use in statistics: Part II. the graphical analysis of the results of a factorial experiment. Applied Statistics, 3(3):184–195, November 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conway:1954:LE

[123] Freda Conway and W. T. Osborn. Letters to the editor. Applied Statistics, 3(3):196–198, November 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:NCc

[124] Anonymous. Notes and comments. Applied Statistics, 3(3):199, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:IASc

[125] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 3(3):200–201, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:SSRc

[126] Anonymous. Study section of the royal statistical society. Applied Statistics, 3(3):202–203, November 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hinsley:1954:BRS

[127] J. F. Hinsley. Book reviews: Statistical Methods in Electrical Engineering, by D. A. Bell. Applied Statistics, 3 (3):204, November 1954. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swan:1954:BRC

[128] A. W. Swan. Book reviews: Le Contrôle Statistique des Fabrications, by R. Cave. Applied Statistics, 3(3):205, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkins:1954:BRS

[129] L. T. Wilkins. Book reviews: Statistical Inference, by H. M. Walker and J. Lev. Applied Statistics, 3(3):206–207, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1954:BRM

[130] D. R. Davies. Book reviews: Mathematics in Action, by O. G. Sutton. Applied Statistics, 3(3):207, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:OPR

[131] Anonymous. Other publications received. Applied Statistics, 3(3):208, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stuart:1954:BRS

[132] A. Stuart. Book reviews: Sampling Theory of Surveys with Applications, by Pandurang V. Sukhatme. Applied Statistics, 3(3):208, November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:BMPc

[133] Anonymous. Back matter. Applied Statistics, 3(3):[1], November 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:FMPc

[134] Anonymous. Front matter. Applied Statistics, 3(3):[2], November 1954. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1954:VIP

[135] Anonymous. Volume information. Applied Statistics, 3(3):[7], November 1954. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beer:1955:PIA

[136] Stafford Beer. The productivity index in active service. Applied Statistics, 4(1): 1–14, March 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Swan:1955:MGS

[137] Albert W. Swan. A method of graphing sales figures for executive use. Applied Statistics, 4(1):15–21, March 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1955:ALP

[138] Noel Williams. An application of linear programming to the selection of raw materials. Applied Statistics, 4(1):22– 31, March 1955. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1955:CMI

[139] Geoffrey H. Jowett. The comparison of means of industrial time series. Applied Statistics, 4(1):32–46, March 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1955:SMM

[140] G. P. Davies and G. W. Sears. Some makeshift methods of analysis applied to complex experimental results. Applied Statistics, 4(1):47–63, March 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sakura:1955:LE

[141] Itasu Sakura. Letter to the editor. Applied Statistics, 4(1):64, March 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vanRest:1955:QA

[142] E. D. van Rest. Question and answer.

Applied Statistics, 4(1):65–66, March 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:IASa

[143] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 4(1):67–69, March 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:SSRa

[144] Anonymous. Study section of the royal statistical society. Applied Statistics, 4 (1):70, March 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Read:1955:BRD

[145] D. R. Read. Book reviews: The Design and Analysis of Industrial Experiments, by O. L. Davies. Applied Statistics, 4 (1):71, March 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1955:BRB

[146] O. L. Davies. Book reviews: Biometrika Tables for Statisticians, Volume I, by E. S. Pearson and H. O. Hartley. Applied Statistics, 4(1):72, March 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:VIPa

[147] Anonymous. Volume information. Applied Statistics, 4(1):[4], March 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jessop:1955:RIP

[148] Marjorie Jessop. Reservoirs of infection of pulmonary tuberculosis in warwickshire. Applied Statistics, 4(2):73– 79, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1955:MRB

[149] Geoffrey H. Jowett. Multiple regression between variables measured at the points of a plane sheet. Applied Statistics, 4(2):80–89, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Daniell:1955:CA

[150] P. A. Daniell and F. D. K. Liddell. Calculation by accumulation. Applied Statistics, 4(2):90–96, June 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Billewicz:1955:SNP

[151] Wladyslaw Z. Billewicz. A simple nonparametric graphical test for the significance of a difference. Applied Statistics, 4(2):97–102, June 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wishart:1955:MA

[152] John Wishart. Multivariate analysis.

Applied Statistics, 4(2):103–116, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Glaskin:1955:LE

[153] A. Glaskin. Letter to the editor. Applied Statistics, 4(2):117–119, June 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:NCa

[154] Anonymous. Notes and comments. Applied Statistics, 4(2):120–121, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:IASb

[155] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 4(2):122–125, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:SSRb

[156] Anonymous. Study section of the royal statistical society. Applied Statistics, 4 (2):126, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1955:BRE

[157] G. H. Jowett. Book reviews: Elements of Statistics, by H. C. Fryer. Applied Statistics, 4(2):127, June 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1955:BRH

[158] L. H. C. Tippett. Book reviews: Handbook of Graphic Presentation, by Calvin F. Schmid. Applied Statistics, 4(2): 127, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moser:1955:BRD

[159] C. A. Moser. Book reviews: The Design of Social Research, by R. L. Ackoff. Applied Statistics, 4(2):128–129, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanRest:1955:BRS

[160] E. D. Van Rest. Book reviews: Statistical Analysis in Chemistry and the Chemical Industry, by Carl A. Bennett and Norman L. Franklin. Applied Statistics, 4(2):129–130, June 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1955:BRQ

[161] L. H. C. Tippett. Book reviews: Quality Control Through Statistical Methods, by Norbert Lloyd Enrick. Applied Statistics, 4(2):130, June 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1955:BRP

[162] L. H. C. Tippett. Book reviews: Productivity, a Critique of Current Usage, by Lewis A. Maverick. Applied Statistics, 4(2):130–131, June 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartlett:1955:BRC

[163] M. S. Bartlett. Book reviews: A concise history of mathematics, by D. J. Struik. Applied Statistics, 4(2):131, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:PRa

[164] Anonymous. Publications received. Applied Statistics, 4(2):132, June 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Saunders:1955:CTN

[165] Christopher T. Saunders. Current trends in national income statistics. Applied Statistics, 4(3):133–144, November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harper:1955:FPS

[166] Roland Harper. Fundamental problems in the subjective appraisal of foodstuffs. Applied Statistics, 4(3):145–161, November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Liston:1955:SFE

[167] David J. Liston. Sales forecasting: An evaluation of some practical results. Applied Statistics, 4(3):162–172, November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hepburn:1955:LQM

[168] John H. Hepburn. Level of quality for minimum cost of manufacture of a specification. Applied Statistics, 4(3): 173–182, November 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shankleman:1955:MRN

[169] Eric Shankleman. Measuring the readership of newspapers and magazines. Applied Statistics, 4(3):183–194, November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ramm:1955:DBS

[170] Bernard T. Ramm. The determination of branch staff establishments in a multiple shop organisation. Applied Statistics, 4(3):195–198, November 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:1955:DTS

[171] Geoffrey M. Clarke. A design for testing several treatments under controlled environmental conditions. Applied Statistics, 4(3):199–206, November 1955. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:NCb

[172] Anonymous. Notes and comments. Applied Statistics, 4(3):207, November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:IASc

[173] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 4(3):208–210, November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:SSRc

[174] Anonymous. Study section of the royal statistical society. Applied Statistics, 4 (3):211, November 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:PRb

[175] Anonymous. Publications received.

Applied Statistics, 4(3):212, November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harland:1955:BRB

[176] S. C. Harland. Book review: Biomathematics: The Principles of Mathematics for Students of Biological Science, by Cedric A. B. Smith. Applied Statistics, 4(3):212, November 1955. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1955:VIPb

[177] Anonymous. Volume information. Applied Statistics, 4(3):[2], November 1955. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goobey:1956:USI

[178] G. H. Ross Goobey. The use of statistics in the investment of funds. Applied Statistics, 5(1):1–11, March 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1956:MCL

[179] David Hill. Modified control limits. Applied Statistics, 5(1):12–19, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandon:1956:RCC

[180] Frank Sandon. A regression control chart for use in personnel selection. Applied Statistics, 5(1):20–31, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harper:1956:FAT

[181] Roland Harper. Factor analysis as a technique for examining complex data on foodstuffs. Applied Statistics, 5(1): 32–48, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cole:1956:FWS

[182] Dorothy E. Cole. Field work in sample surveys of household income and expenditure. Applied Statistics, 5(1):49– 61, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:NCa

[183] Anonymous. Notes and comments. Applied Statistics, 5(1):62, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Plackett:1956:QA

[184] R. L. Plackett. Questions and answers. Applied Statistics, 5(1):63–64, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:IASa

[185] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 5(1):65–66, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:SSRa

[186] Anonymous. Study section of the royal statistical society. Applied Statistics, 5 (1):67, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Plackett:1956:BRC

[187] R. L. Plackett. Book reviews: Chance and Choice by Cardpack and Chessboard. Volume 2, by Lancelot Hogben. Applied Statistics, 5(1):68, March 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1956:BRS

[188] G. H. Jowett. Book reviews: Statistical Methods for Research Workers, by Ronald A. Fisher. Applied Statistics, 5 (1):68–70, March 1956. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1956:BRI

[189] G. H. Jowett. Book reviews: An Introduction to Stochastic Processes, by M. S. Bartlett. Applied Statistics, 5(1): 70, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shankleman:1956:BRP

[190] E. Shankleman. Book reviews: Population Statistics and their Compilation, by Hugh H. Wolfenden. Applied Statistics, 5(1):71–72, March 1956. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:PRa

[191] Anonymous. Publications received. Applied Statistics, 5(1):72, March 1956.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:BMPa

[192] Anonymous. Back matter. Applied Statistics, 5(1):[2], March 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:FMPa

[193] Anonymous. Front matter. Applied Statistics, 5(1):[2], March 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gray:1956:EIV

[194] Percy G. Gray. Examples of interviewer variability taken from two sample surveys. Applied Statistics, 5(2): 73–85, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

James:1956:SSP

[195] Stanley F. James. Some sampling problems in connection with accounting records. Applied Statistics, 5(2): 86–105, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gridgeman:1956:TE

[196] Norman Theodore Gridgeman. A tasting experiment. Applied Statistics, 5(2): 106–112, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ferber:1956:CGP

[197] Robert Ferber. Are correlations any guide to predictive value? Applied Statistics, 5(2):113–121, June 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnett:1956:FEF

[198] Martin K. Barnett and Frank C. Mead, Jr. A 2 4 factorial experiment in four blocks of eight: A study in radioactive decontamination. Applied Statistics, 5 (2):122–131, June 1956. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Leander:1956:EUT

[199] Erik K. Leander and David J. Finney.

An extension of the use of the χ^{2}-test. Applied Statistics, 5(2):132–136, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1956:QA

[200] G. H. Jowett. Question and answer.

Applied Statistics, 5(2):137–138, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:IASb

[201] Anonymous. Industrial applications section of the royal statistical society. Applied Statistics, 5(2):139, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:SSRb

[202] Anonymous. Study section of the royal statistical society. Applied Statistics, 5 (2):140, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartlett:1956:BRS

[203] M. S. Bartlett. Book reviews: Stochastic Models for Learning, by R. R. Bush and F. Mosteller. Applied Statistics, 5 (2):141, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Utting:1956:BRC

[204] J. E. G. Utting. Book reviews: Contributions of Survey Methods to Economics, by George Katona, Lawrence R. Klein, John B. Lansing and James N. Morgan. Applied Statistics, 5(2):141–143, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkins:1956:BRC

[205] L. T. Wilkins. Book reviews: The Communication of Ideas, by T. Cauter and J. S. Downham. Applied Statistics, 5(2): 143–144, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1956:BRH

[206] D. G. Beech. Book reviews: Handbook of Industrial Engineering and Management, by W. G. Ireson and E. L. Grant. Applied Statistics, 5(2):144, June 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:BMPb

[207] Anonymous. Back matter. Applied Statistics, 5(2):[1], June 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:FMPb

[208] Anonymous. Front matter. Applied Statistics, 5(2):[2], June 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1956:SHP

[209] Norman T. J. Bailey. Statistics in hospital planning and design. Applied Statistics, 5(3):146–157, November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jessop:1956:MCM

[210] W. Neil Jessop. Monte Carlo methods and industrial problems. Applied Statistics, 5(3):158–165, November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1956:ASM

[211] Noel Williams. The allocation of scarce materials between products. Applied Statistics, 5(3):166–176, November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maurice:1956:ICG

[212] Rita J. Maurice. Industrial censuses in Great Britain and the United States. Applied Statistics, 5(3):177–194, November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Belson:1956:TSE

[213] William A. Belson. A technique for studying the effects of a television broadcast. Applied Statistics, 5(3):195–202, November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1956:MVE

[214] Michael Healy and Michael Westmacott.

Missing values in experiments analysed on automatic computers. Applied Statistics, 5(3):203–206, November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Banerjee:1956:CCP

[215] K. S. Banerjee. A comment on the construction of price index numbers. Applied Statistics, 5(3):207–210, November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:NCb

[216] Anonymous. Notes and comments. Applied Statistics, 5(3):211–212, November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:MSR

[217] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 5(3):213–214, November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1956:BRP

[218] Norman T. J. Bailey. Book reviews: Prediction Methods in Relation to Borstal Training, by Hermann Mannheim and Leslie T. Wilkins. Applied Statistics, 5(3):215–216, November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:PRb

[219] Anonymous. Publications received. Applied Statistics, 5(3):216, November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1956:BRC

[220] D. G. Beech. Book reviews: Correlated Random Normal Deviates, by E. C. Fieller, T. Lewis and E. S. Pearson.

Applied Statistics, 5(3):216, November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:BMPc

[221] Anonymous. Back matter. Applied Statistics, 5(3):[2], November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:FMPc

[222] Anonymous. Front matter. Applied Statistics, 5(3):[2], November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:FP

[223] Anonymous. Foreword. Applied Statistics, 5(3):[1], November 1956. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1956:VIP

[224] Anonymous. Volume information. Applied Statistics, 5(3):[5], November 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Klein:1957:SLE

[225] L. R. Klein. The scope and limitations of econometrics. Applied Statistics, 6(1): 1–17, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cartwright:1957:EOI

[226] Ann Cartwright. The effect of obtaining information from different informats on a family morbidity inquiry. Applied Statistics, 6(1):18–25, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kay:1957:LPP

[227] Emil Kay and Eric Duckworth. Linear programming in practice. Applied Statistics, 6(1):26–39, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tulse:1957:SVV

[228] R. Tulse. Sampling for variables with a very skew distribution. Applied Statistics, 6(1):40–44, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baker:1957:APR

[229] Anthony G. Baker. Analysis and presentation of the results of factorial experiments. Applied Statistics, 6(1):45– 55, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chittleburgh:1957:PPC

[230] Raymond Chittleburgh and Alec Lee. A problem in production control. Applied Statistics, 6(1):56–62, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stannage:1957:URA

[231] William Stannage. Use of regression analysis to detect errors in measurement of intermediate materials in a multistage process. Applied Statistics, 6(1): 63–66, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brayer:1957:CSE

[232] Edward F. Brayer. Calculating the standard error of a proportion. Applied Statistics, 6(1):67–68, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:NCa

[233] Anonymous. Notes and comments. Applied Statistics, 6(1):69–70, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:MSRa

[234] Anonymous. Meetings of sections of the royal statistical societ. Applied Statistics, 6(1):71–72, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1957:BRE

[235] Cedric A. B. Smith. Book reviews: Experimental Design and Its Statistical Basis, by D. J. Finney. Applied Statistics, 6 (1):73, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Saunders:1957:BRG

[236] C. T. Saunders. Book reviews: Guide to New Zealand Official Statistics, by E. P. Neale. Applied Statistics, 6(1):74, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1957:BRC

[237] D. G. Beech. Book reviews: Curso de Estadistica: Volume I, Estadistica Descriptiva y Modelo Matematico, by Enrique Chacon. Applied Statistics, 6(1): 74–75, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1957:BRN

[238] D. G. Beech. Book reviews: Nellie Landblom’s Copybook for Beginners in Research Work, by Nellie T. Landblom. Applied Statistics, 6(1):75, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkins:1957:BRP

[239] Leslie T. Wilkins. Book reviews: A primer of social statistics, by S. M. Dornbusch and C. F. Schmid. Applied Statistics, 6(1):75–76, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1957:BRD

[240] L. H. C. Tippett. Book reviews: Dynamic Factors in Industrial Productivity, by Seymour Melman. Applied Statistics, 6(1):76–77, March 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nuttall:1957:BRC

[241] C. Nuttall. Book reviews: The Classification and Coding of Accounts, by J. M. S. Risk. Applied Statistics, 6(1): 78, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandon:1957:BRE

[242] Frank Sandon. Book reviews: The Essentials of Educational Statistics, by Francis G. Cornell. Applied Statistics, 6(1):78–79, March 1957. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coutie:1957:BRT

[243] G. A. Coutie. Book reviews: The Theory of Games and Linear Programming, by S. Vajda. Applied Statistics, 6(1): 79–80, March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:VIPa

[244] Anonymous. Volume information. Applied Statistics, 6(1):[4], March 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Box:1957:EOM

[245] George E. P. Box. Evolutionary operation: A method for increasing industrial productivity. Applied Statistics, 6 (2):81–101, June 1957. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lipton:1957:SSA

[246] Stephen Lipton. Some statistical applications of electronic computers. Applied Statistics, 6(2):102–113, June 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gibson:1957:TGRa

[247] Wendy M. Gibson and Geoffrey H.

Jowett. ‘three-group’ regression analysis: Part I. simple regression analysis. Applied Statistics, 6(2):114–122, June 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gales:1957:DFS

[248] Kathleen Gales. Discriminant functions of socio-economic class. Applied Statistics, 6(2):123–132, June 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morrison:1957:SVE

[249] S. James Morrison. The study of variability in engineering design. Applied Statistics, 6(2):133–138, June 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gray:1957:SSB

[250] Percy G. Gray. A sample survey with both a postal and an interview stage. Applied Statistics, 6(2):139–153, June 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:NCb

[251] Anonymous. Notes and comments. Applied Statistics, 6(2):154, June 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:MSRb

[252] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 6(2):155–156, June 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1957:BRS

[253] L. H. C. Tippett. Book reviews: Statistics: A New Approach, by W. Allen Wallis and Harry V. Roberts. Applied Statistics, 6(2):157–158, June 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Burke:1957:BRS

[254] S. Burke. Book reviews: Statistics and Their Application to Commerce. 11th Edition, by A. R. Ilersic. Applied Statistics, 6(2):158–159, June 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:PRa

[255] Anonymous. Publications received. Applied Statistics, 6(2):160, June 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1957:BRG

[256] D. G. Beech. Book reviews: Guide to Elementary Statistical Formulas, by Robert E. Johnson and Doris N. Morris. Applied Statistics, 6(2):160, June 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tanner:1957:SRT

[257] John C. Tanner. The sampling of road traffic. Applied Statistics, 6(3):161–170, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moore:1957:NQC

[258] Peter G. Moore. Normality in quality control charts. Applied Statistics, 6(3): 171–179, November 1957. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barr:1957:DBE

[259] Alexander Barr. Differences between experienced interviewers. Applied Statistics, 6(3):180–188, November 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gibson:1957:TGRb

[260] Wendy M. Gibson and Geoffrey H.

Jowett. ‘three-group’ regression analysis: Part II. multiple regression analysis. Applied Statistics, 6(3):189–197, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tomlinson:1957:SSP

[261] Rolfe C. Tomlinson. A simple sequential procedure to test whether average conditions achieve a certain standard. Applied Statistics, 6(3):198–207, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemp:1957:OPU

[262] Kenneth W. Kemp. An optical periodograph for use in the analysis of continuous measurements. Applied Statistics, 6(3):208–213, November 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stevens:1957:CLR

[263] Colin Stevens. Confidence limits for the range of a search radar. Applied Statistics, 6(3):214–222, November 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:NCc

[264] Anonymous. Notes and comments. Applied Statistics, 6(3):223, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:MSRc

[265] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 6(3):224, November 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Saunders:1957:BRN

[266] C. T. Saunders. Book reviews: National Income Visualized, by Arthur O. Dahlberg. Applied Statistics, 6(3):225– 226, November 1957. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1957:BRS

[267] G. H. Jowett. Book reviews: Statistical Methods and Scientific Inference, by Ronald A. Fisher. Applied Statistics, 6(3):226–227, November 1957. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oldham:1957:BRD

[268] P. D. Oldham. Book reviews: A dictionary of statistical terms, by Maurice G. Kendall and William R. Buckland. Applied Statistics, 6(3):227–228, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1957:BRL

[269] W. J. Corlett. Book reviews: The Lognormal Distribution, With Special Reference to Its Uses in Economics, by J. Aitchison and J. A. C. Brown. Applied Statistics, 6(3):228–230, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1957:BRS

[270] D. G. Beech. Book reviews: Statistical Methods in Research and Production, With Special Reference to the Chemical Industry. Third Edition, by O. L. Davies. Applied Statistics, 6(3):230–231, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rowntree:1957:BRP

[271] J. A. Rowntree. Book reviews: The 1953 Pilot Population Census for the First Population Census in Sudan. Applied Statistics, 6(3):231–232, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:PRb

[272] Anonymous. Publication received. Applied Statistics, 6(3):232, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1957:VIPb

[273] Anonymous. Volume information. Applied Statistics, 6(3):[3], November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Armitage:1958:SVT

[274] Peter Armitage, G. Isley, and C. Ingram. Studies in the variability of two bloodclotting test. Applied Statistics, 7(1): 1–14, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maurice:1958:OSE

[275] Rita J. Maurice. Official statistics of economic trends. Applied Statistics, 7(1): 15–24, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Friend:1958:PAO

[276] John K. Friend. Probits: An application to observational data. Applied Statistics, 7(1):25–41, March 1958. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Duncan:1958:SAS

[277] David C. Duncan. Statistics of an auction sale. Applied Statistics, 7(1):42– 47, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mansfield:1958:RCC

[278] Edwin Mansfield and Harold H. Wein. A regression control chart for cost. Applied Statistics, 7(1):48–57, March 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gray:1958:ILS

[279] Percy G. Gray. Initial letters of surnames. Applied Statistics, 7(1):58– 59, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:NCa

[280] Anonymous. Notes and comments. Applied Statistics, 7(1):60–61, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:MSRa

[281] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 7(1):62–64, March 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beer:1958:BRRa

[282] Stafford Beer. Book reviews: The Rationalization of Industrial Decision Processes, by William T. Morris. Applied Statistics, 7(1):65–66, March 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beer:1958:BRRb

[283] Stafford Beer. Book reviews: The Reduction and Presentation of Experimental Results, by J. T. Richardson. Applied Statistics, 7(1):66–67, March 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanRest:1958:BRP

[284] E. D. Van Rest. Book reviews: Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol. V. Econometrics, Industrial Research, and Psychometry, by J. Neyman. Applied Statistics, 7(1):67– 68, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walker:1958:BRS

[285] A. M. Walker. Book reviews: Statistical Analysis of Stationary Time Series, by U. Grenander and M. Rosenblatt. Applied Statistics, 7(1):68–70, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1958:BRF

[286] D. G. Beech. Book reviews: The Fascination of Numbers, by W. J. Reichmann. Applied Statistics, 7(1):70, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1958:BRN

[287] F. T. M. Smith. Book reviews: Numerical Methods, by R. A. Buckingham. Applied Statistics, 7(1):70–71, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anson:1958:BRI

[288] C. J. Anson. Book reviews: Introduction to Statistical Inference, by Jerome C. R. Li. Applied Statistics, 7(1):71–72, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:PRa

[289] Anonymous. Publications received. Applied Statistics, 7(1):72, March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:VIPa

[290] Anonymous. Volume information. Applied Statistics, 7(1):[4], March 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:E

[291] Anonymous. Editorial. Applied Statistics, 7(2):73–75, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shankleman:1958:EHS

[292] Eric Shankleman. The economist and his statistics. Applied Statistics, 7(2): 76–85, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Broadbent:1958:SMR

[293] Simon Broadbent. Simple mortality rates. Applied Statistics, 7(2):86– 95, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Muir:1958:MC

[294] Andrew Muir. Marginal cost. Applied Statistics, 7(2):96–103, June 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1958:SAM

[295] Norman T. J. Bailey. A simple approach to the minimisation of store-size. Applied Statistics, 7(2):104–113, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1958:FA

[296] Geoffrey H. Jowett. Factor analysis. Applied Statistics, 7(2):114–125, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:NCb

[297] Anonymous. Notes and comments. Applied Statistics, 7(2):126, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:MSRb

[298] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 7(2):127–129, June 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nuttall:1958:BRS

[299] C. Nuttall. Book reviews: Statistical Sources for Market Research. Applied Statistics, 7(2):130, June 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stanley:1958:BRI

[300] Philip J. Stanley. Book reviews: Introduction to Statistical Reasoning, by Philip J. McCarthy. Applied Statistics, 7(2):130–131, June 1958. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:PRb

[301] Anonymous. Publications received. Applied Statistics, 7(2):132, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jessop:1958:BRR

[302] W. N. Jessop. Book reviews: Regression Analysis of Production Costs and Factory Operations. (Third Edition), by Philip Lyle and Leonard H. C. Tippett. Applied Statistics, 7(2):132, June 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tippett:1958:GAS

[303] Leonard H. C. Tippett. A guide to acceptance sampling. Applied Statistics, 7(3):133–148, November 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hamaker:1958:SBP

[304] Hugo C. Hamaker. Some basic principles of sampling inspection by attributes. Applied Statistics, 7(3):149– 159, November 1958. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morrison:1958:LDQ

[305] James Morrison. The lognormal distribution in quality control. Applied Statistics, 7(3):160–172, November 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Glenn:1958:AVR

[306] William Alexander Glenn and Clyde Young Kramer. Analysis of variance of a randomised block design with missing observations. Applied Statistics, 7(3):173– 185, November 1958. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lindley:1958:PHC

[307] D. V. Lindley. Professor Hogben’s ‘crisis’ — A survey of the foundations of statistics. Applied Statistics, 7(3):186–198, November 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:NCc

[308] Anonymous. Notes and comments. Applied Statistics, 7(3):199, November 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:MSRc

[309] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 7(3):200, November 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barr:1958:BRS

[310] A. Barr. Book reviews: Social Aspects of Prescribing, by J. P. Martin. Applied Statistics, 7(3):201–202, November 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

VanRest:1958:BRS

[311] E. D. Van Rest. Book reviews: Statistical Methods in Quality Control, by Dudley J. Cowden. Applied Statistics, 7(3): 202–203, November 1958. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Foster:1958:BRI

[312] F. G. Foster. Book reviews: An Introduction to Probability Theory and its Applications, Volume 1 (2nd Edition), by William Feller. Applied Statistics, 7(3):203–204, November 1958. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:PRc

[313] Anonymous. Publications received.

Applied Statistics, 7(3):204, November 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1958:VIPb

[314] Anonymous. Volume information. Applied Statistics, 7(3):[3], November 1958. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Newbery:1959:SAQ

[315] Geoffrey Newbery. Some aspects of quality control in cotton textiles. Applied Statistics, 8(1):1–12, March 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1959:USE

[316] G. H. Freeman. The use of the same experimental material for more than one set of treatments. Applied Statistics, 8 (1):13–20, March 1959. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1959:EBP

[317] Denis H. Ward. Elimination of bias in the preliminary examination of data. Applied Statistics, 8(1):21–25, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1959:PCP

[318] A. S. C. Ehrenberg. The pattern of consumer purchases. Applied Statistics, 8 (1):26–41, March 1959. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Banerjee:1959:CCF

[319] K. S. Banerjee. Choice of conversion factor in the derivation of an index of a discontinued series. Applied Statistics, 8(1): 42–44, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosenbaum:1959:SCP

[320] Sidney Rosenbaum. A significance chart for percentages. Applied Statistics, 8(1): 45–52, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:NCa

[321] Anonymous. Notes and comments. Applied Statistics, 8(1):53–54, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:MSRa

[322] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 8(1):55–57, March 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Emmett:1959:BRM

[323] B. P. Emmett. Book reviews: Motivation Research, by Harry Henry. Applied Statistics, 8(1):58, March 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1959:BRM

[324] D. G. Beech. Book reviews: Multivariate Correlation Analysis, by Philip H. DuBois. Applied Statistics, 8(1):58– 60, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oldham:1959:BRM

[325] P. D. Oldham. Book reviews: The Mathematical Theory of Epidemics, by N. T. J. Bailey. Applied Statistics, 8(1):60– 61, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1959:BRAa

[326] D. G. Beech. Book reviews: The Advanced Theory of Statistics. Volume 1, Distribution Theory, by Maurice G. Kendall and Alan Stuart. Applied Statistics, 8(1):61–62, March 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Emmett:1959:BRS

[327] B. P. Emmett. Book reviews: Sampling Opinions: An Analysis of Survey Procedure, by Frederick F. Stephan and Philip J. McCarthy. Applied Statistics, 8(1): 62–63, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jackson:1959:BRQ

[328] R. R. P. Jackson. Book reviews:

Queues, Inventories and Maintenance: The Analysis of Operational Systems with Variable Demand and Supply, by Philip M. Morse. Applied Statistics, 8 (1):63, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:PRa

[329] Anonymous. Publications received. Applied Statistics, 8(1):64, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bulmer:1959:BRF

[330] M. G. Bulmer. Book reviews: Fundamentals of Statistical Reasoning, by M. H. Quenouille. Applied Statistics, 8(1): 64, March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:VIPa

[331] Anonymous. Volume information. Applied Statistics, 8(1):[4], March 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Belson:1959:MPP

[332] William A. Belson. Matching and prediction on the principle of biological classification. Applied Statistics, 8(2): 65–75, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1959:DSS

[333] I. D. Hill, Gareth Horsnell, and Bernard T. Warner. Deferred sentencing schemes. Applied Statistics, 8(2): 76–91, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gray:1959:LCS

[334] Percy G. Gray and Elizabeth A. Parr. The length of cigarette stubs. Applied Statistics, 8(2):92–103, June 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1959:UEP

[335] Denis H. Ward. The use of edge-punched cards in statistical computation. Applied Statistics, 8(2):104–113, June 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemsley:1959:DBS

[336] W. F. F. Kemsley. Designing a budget survey. Applied Statistics, 8(2):114– 123, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Page:1959:PRE

[337] E. S. Page. Pseudo-random elements for computers. Applied Statistics, 8(2): 124–131, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hil:1959:LE

[338] I. D. Hil. Letter to the editor. Applied Statistics, 8(2):132, June 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:MSRb

[339] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 8(2):133–134, June 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1959:BRA

[340] G. H. Jowett. Book reviews: The Analysis of Multiple Time-Series, by M. H. Quenouille. Applied Statistics, 8(2): 135, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkins:1959:BRC

[341] L. T. Wilkins. Book reviews: A course in multivariate analysis, by Maurice G. Kendall. Applied Statistics, 8(2):135– 136, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandon:1959:BRF

[342] Frank Sandon. Book reviews: A first course in statistics, by R. Loveday. Applied Statistics, 8(2):136–137, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sandon:1959:BRP

[343] Frank Sandon. Book reviews: Principles of Statistical Techniques, by P. G. Moore. Applied Statistics, 8(2):136– 137, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkins:1959:BRS

[344] L. T. Wilkins. Book reviews: Survey Methods in Social Investigation, by Claus A. Moser. Applied Statistics, 8(2): 137–139, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1959:BRE

[345] W. J. Corlett. Book reviews: Economic Arithmetic, by Robin Marris. Applied Statistics, 8(2):139–142, June 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lucas:1959:BRO

[346] P. G. Lucas. Book reviews: Operational Research in Practice, by Max Davies and Michel Verhulst. Applied Statistics, 8 (2):142, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Corlett:1959:BRA

[347] W. J. Corlett. Book reviews: Applied Statistics for Economists, by P. H. Karmel. Applied Statistics, 8(2):142– 143, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1959:BRAb

[348] D. G. Beech. Book reviews: Application of Statistics to Paper Testing. Applied Statistics, 8(2):143–144, June 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:PRb

[349] Anonymous. Publications received. Applied Statistics, 8(2):144, June 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

James:1959:SAS

[350] John S. James. A statistical approach to stores auditing. Applied Statistics, 8(3): 145–157, November 1959. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ma:1959:ASQ

[351] Ronald Ma. The age structure of quoted public companies. Applied Statistics, 8(3):158–167, November 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ashford:1959:PSC

[352] John R. Ashford. A problem of subjective classification in industrial medicine. Applied Statistics, 8(3):168– 185, November 1959. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1959:RSS

[353] A. S. C. Ehrenberg and J. M. Shewan. A reliability study of sensory assessments. I. Applied Statistics, 8(3): 186–195, November 1959. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chisholm:1959:SCT

[354] Michael Chisholm. Shipping costs and the terms of trade: Australia and new zealand. Applied Statistics, 8(3):196– 201, November 1959. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ward:1959:WMQ

[355] Denis H. Ward. Weekly, monthly, and quarterly tolerances for coke quality. Applied Statistics, 8(3):202–209, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hamaker:1959:ASS

[356] Hugo C. Hamaker. Adjusting single sampling plans for finite lot size. Applied Statistics, 8(3):210–214, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:NCb

[357] Anonymous. Notes and comments. Applied Statistics, 8(3):215–216, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:MSRc

[358] Anonymous. Meetings of sections of the royal statistical society. Applied Statistics, 8(3):217–218, November 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mackenzie:1959:BRI

[359] H. C. Mackenzie. Book reviews: Introduction to Agricultural Economic Analysis, by C. E. Bishop and W. D. Toussaint. Applied Statistics, 8(3):219–220, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stein:1959:BRR

[360] Lilli Stein. Book reviews: Recent Studies in Epidemiology, by J. Pemberton and H. Willard. Applied Statistics, 8(3): 220–221, November 1959. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beer:1959:BRW

[361] Stafford Beer. Book reviews: What Is Cybernetics?, by G. T. Guilbaud. Applied Statistics, 8(3):221, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beech:1959:BRP

[362] D. G. Beech. Book reviews: Planning of Experiments, by D. R. Cox. Applied Statistics, 8(3):222, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Emmett:1959:BRT

[363] B. P. Emmett. Book reviews: Television and the Child, by Hilde T. Himmelweit, A. N. Oppenheim and Pamela Vince. Applied Statistics, 8(3):222–223, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ridley:1959:BRB

[364] T. M. Ridley. Book reviews: Business Budgets and Accounts, by Harold C. Edey. Applied Statistics, 8(3):224, November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:BMP

[365] Anonymous. Back matter. Applied Statistics, 8(3):[1], November 1959. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1959:VIPb

[366] Anonymous. Volume information. Applied Statistics, 8(3):[2], November 1959. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).